# Reasons to Reject? Aligning Language Models with Judgments

### v1

@@ 4.1 Incorporating Judgments for Alignment
We call an instruction-response pair “aligned” if the response follows the instruction faithfully and satisfies human expectations $\bm{x}\xrightarrow{}\bm{y}$. Otherwise, a judgment describes the errors or deficiencies present in the response. Assuming the task is to generate a response that intentionally fulfills the judgment, it can be inferred that the response always aligns with the combined input of instruction and judgment $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$. Based on the idea, we construct three types of alignment data, depicted in Table 1.
Align-P:
The LLM produces a satisfactory response $\bm{y}$ to the original instruction $\bm{x}$. Therefore, a positive judgment $\bm{j}$ is conferred to acknowledge the commendable performance of the LLM. It is evident that the response $\bm{y}$ is aligned with the instruction $\bm{x}$ as well as the combined input $[\bm{x},\bm{j}]$.
Align-N:
The LLM makes some mistakes in its generation, resulting in an unsatisfactory response $\bm{y}$. Consequently, a negative judgment $\bm{j}$ details the corresponding critiques. For Align-N, $\bm{y}$ is not aligned with original instruction $\bm{x}$. However, when considering $\bm{x}$ and $\bm{j}$ as a whole, $\bm{y}$ is indeed aligned with the combined input $[\bm{x},\bm{j}]$.
Misalign:
The authentic negative judgment in Align-N is substituted with a fake positive judgment $\bm{j}$. In this case, the response $\bm{y}$ is not aligned with either the original instruction $\bm{x}$ or the combination of instruction and judgment $[\bm{x},\bm{j}]$.
@@ 4.2 Learning from Contrasting
Figure 2: The probability of generating identical output text in an LLM under Align-N (left) and Misalign (right) contexts.
Align-N vs. Misalign:
Despite Align-N and Misalign are not aligned in terms of $\bm{x}\xrightarrow{}\bm{y}$, they show opposite polarities in the task of $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$. Thanks to the strong in-context learning capabilities of LLMs, the alignment flip from Align-N (aligned) to Misalign (misaligned) is often accompanied by decreased generation probabilities of the response, particularly for tokens that exhibit a strong correlation with the authentic negative judgment. Figure 2 presents a simple example, wherein the response commits a minor capitalization issue. The LLM assigns a considerably higher probability for “a” when taking the authentic negative judgment $\bm{j}^{-}$ instead of the fake positive judgment $\bm{j}^{+}$ as additional input, precisely at the point where the LLM commits the error.
To take advantage of the above contrast, we feed Align-N and Misalign examples to the LLM to get token generation probabilities $p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})$ and $p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{+})$ separately. We consider the tokens that display a substantially increased generation probability when conditioned on $\bm{j}^{-}$ compared to $\bm{j}^{+}$ as inappropriate tokens (e.g., “a” in Figure 2). Concretely, the following criterion is adopted:
$U=\{t\;|\;p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})-\lambda\cdot p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{+})>0\}$ (5)
where $\lambda\geq 1$ is a hyperparameter to tradeoff the precision and recall of detecting inappropriate tokens.
We apply the UT objective (Welleck et al., 2020) on the identified inappropriate tokens for pushing the LLM to explore alternative generations. For other tokens, we use the standard MLE loss:
$L_{1}=-\frac{1}{N}(\sum_{t\notin U}\log p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})+\sum_{t\in U}\alpha\log(1-p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})))$ (6)
where $\alpha$ is another hyperparameter to control the scale of unlikelihood loss.
Align-P vs. Align-N:
Despite both Align-P and Align-N are aligned in terms of $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$, only Align-P is aligned when solely considering the instruction ($\bm{x}\xrightarrow{}\bm{y}$). Essentially, it suggests that the LLM should output different responses depending on whether a negative judgment is incorporated or not. Therefore, the comparison provides valuable information for the LLM to discern satisfactory and unsatisfactory responses. Specifically, we train on this comparison with the following MLE objective:
$L_{2}=-\frac{\mathbbm{1}(\bm{x}\xrightarrow{}\bm{y})}{N}\sum_{t}\log p_{\theta}(y_{t}|\bm{y}_{<t},\bm{x})-\frac{(1-\mathbbm{1}(\bm{x}\xrightarrow{}\bm{y}))}{N}\sum_{t}\log p_{\theta}(y_{t}|\bm{y}_{<t},\bm{j},\bm{x})$ (7)
@@
Finally, the overall loss of CUT combines the loss functions in the two contrasts: $L_{\text{CUT}}=L_{1}+L_{2}$.
@@ 4.3 Relation to Prior Solutions
  • Forward Prediction: Forward prediction teaches an LLM to generate judgments in the hope of indirectly boosting its response generation abilities. In contrast, CUT directly utilizes judgments to teach the LLM how to generate satisfactory responses and avoid unsatisfactory ones.
  • ILF: ILF assumes judgments can elicit improved responses and solely learn from such pseudo-aligned instruction-response pairs. Conversely, CUT can directly learn from misaligned instruction-response pairs.
  • Hindsight: Hindsight learns to generate responses of different qualities at the risk of increasing the likelihood of unsatisfactory responses. In comparison to Hindsight, CUT mitigates this issue by incorporating both likelihood and unlikelihood training objectives, as well as inappropriate token detection.
@@ 5 Experiments
We experiment with CUT in two alignment settings: (1) Offline alignment where off-the-shelf model-agnostic instruction-response-judgment triplets are used. (2) Online alignment where the judgments are made based on responses generated by the current target model. This online setting can be implemented iteratively, allowing for continuous refinement and adaptation.
Implementations.
We train our models using LoRA (Hu et al., 2022) and follow the best configurations suggested by Platypus (Lee et al., 2023). The tradeoff hyperparameter $\lambda$ is selected from $\{1.1,1.2,1.5\}$ and the unlikelihood weight $\alpha$ is selected from $\{0.25,0.5,0.75,1\}$. We adopt the Alpaca template (Taori et al., 2023) for fine-tuning and inference. The details are in Appendix A.1.
@@ 5.1 Offline Alignment
The offline setting utilizes off-the-shelf instruction-response-judgment triplets for alignment. This aims to check the feasibility of the CUT framework in learning from judgments prior to initiating the costly process of model-specific judgment annotation.
Tasks.
  • General Instruction-following: We train models on the Shepherd dataset (Wang et al., 2023a), which consists of judgment data on diverse NLP tasks such as math word problems and commonsense reasoning. There are 1317 examples in total. For evaluation, we report model performance on four ranking-based and one generation-based LLM benchmarks [^4] Ranking-based evaluation tests an LLM’s ability to select the best response from a set of candidate responses, while generation-based evaluation assesses an LLM’s ability to generate high-quality responses.. Following the Open LLM Leaderboard (Gao et al., 2021), the ranking-based benchmarks are 25-shot ARC (Clark et al., 2018), 10-shot HellaSwag (Zellers et al., 2019), 5-shot MMLU (Hendrycks et al., 2021), and 0-shot TruthfulQA (Lin et al., 2022). The generation-based benchmark is AlpacaEval [^5] Following conventions, GPT4 is utilized to judge the winning rate of the responses generated by our models against those produced by DaVinci003..
Setup.
We use two different base models, LLaMA2-13b and LLaMA2-chat-13b, aiming to demonstrate the efficacy of CUT in both cold-start (LLaMA2) and warm-start (LLaMA2-chat) scenarios. The baseline methods include the base model without further fine-tuning, and the three methods for aligning LLMs with judgments, as discussed in Section 3.2: ILF, Forward Prediction, and Hindsight. Additionally, we establish a baseline, namely Demonstration, in which we fine-tune the base LLM using only instruction-response pairs while ignoring the associated judgments. The comparison between judgment-engaged baselines and Demonstration can better shed light on the benefits of utilizing judgments.
Model Judgment Objective ARC HellaSwag MMLU TruthfulQA Avg. AlpacaEval
[…]
Table 2: Results on the general instruction-following task. The Judgment column indicates if the method utilizes judgments during the alignment. The Objective column denotes the training objective of the alignment stage.
Model Judgment Objective rouge1 rouge2 rougeL rougeLsum
[…]
Table 3: Results on the summarization task.
Results.
@@
The results of the general instruction-following and summarization tasks are presented in Table 2 and Table 3, respectively. For the cold-start scenarios where LLaMA2 is used as the base model, CUT improves the winning rate on AlpacaEval from 1.87 to 62.56 and surpasses the best baseline (i.e., Hindsight) by 52.34 points. This achievement is particularly noteworthy considering that the resulting 13B model, which is fine-tuned with merely 1317 examples, can also beat the 175B DaVinci003. Moreover, CUT improves the base model by 13.08 points on TruthfulQA. This observation implies that CUT can effectively mitigate the issue of hallucination. Conversely, most baselines experience considerable performance deterioration on TruthfulQA. This is likely due to their application of the MLE objective on error-prone responses, which consequently results in diminished factuality in response generation. In terms of ARC, HellaSwag, and MMLU, CUT’s performance remains competitive with the base model, indicating CUT suffers less from the alignment tax problem (Ouyang et al., 2022). For our single NLP task (i.e., summarization) experiments, CUT also surpasses the best baseline (i.e., Forward Prediction) by 1.21 rougeLsum scores. Overall, the results show that CUT is effective in transforming LLMs into both performant generalist and specialist models. Conversely, the performance of other judgment-engaged baselines does not exhibit any notable differences when compared to Demonstration across the five evaluated benchmarks. These results suggest that prior methods cannot effectively utilize judgments.
@@
For the warm-start scenarios where LLaMA2-chat is used as the base model, the performance improvements are consistent with the cold-start scenarios, demonstrating the effectiveness of CUT in learning from judgments in both cold-start and warm-start scenarios. Interestingly, ILF outperforms Forward Prediction and Hindsight on AlpacaEval in warm-start scenarios, despite that the opposite outcome is observed in cold-start scenarios. This may be attributed to that ILF heavily relies on the base model in producing high-quality improved responses, making it less effective in cold-start scenarios.
Model Instruction-following Summarization
@@
CUT 48.84 40.05
 - $L_{1}$ 39.01 37.46
 - first part of $L_{2}$ 27.73
 - second part of $L_{2}$ 46.42 33.60
 - Inappropriate Token Detection 0 0
Table 4: Ablation study on the design of CUT. We report the results on TruthfulQA (Acc.) and summarization test set (rougeLsum) for the two tasks respectively. “-” means that the training set of general instruction-following does not contain Align-P examples. “0” indicates that the training fails to converge.
Ablation Study.
To investigate the effectiveness of the two contrasts employed by CUT, we further perform ablation experiments by eliminating certain training signals. The results are shown in Table 4. We can see that, upon removing the contrast between Align-N and Misalign (- $L_{1}$), the performance of TruthfulQA substantially declines. This finding highlights that the UT objective plays a crucial role in mitigating the hallucination issue. The exclusion of the contrast between Align-P and Align-N can be implemented in two ways. We can either remove the first part or the second part of $L_{2}$. As seen, the impact of removing Align-P is more pronounced than removing Align-N on the summarization task. This may be attributed to the necessity of positive examples for adapting the LLM to a specific task. Furthermore, we introduce an additional ablated variant in which the inappropriate token detection process (Eq. 5) is omitted (- Inappropriate Token Detection). Concretely, we simply apply UT for all tokens in misaligned responses instead. Intriguingly, we find that this approach fails to converge during training. This observation underscores the importance of inappropriate token detection.

### v2

@@ 4.1 Incorporating Judgments for Alignment
We call an instruction-response pair “aligned" if the response follows the instruction faithfully and satisfies human expectations $\bm{x}\xrightarrow{}\bm{y}$. Otherwise, a judgment describes the errors or deficiencies present in the response. Assuming the task is to generate a response that intentionally fulfills the judgment, it can be inferred that the response always aligns with the combined input of instruction and judgment $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$. Based on the idea, we construct three types of alignment data, depicted in Table 1.
Align-P: The LLM produces a satisfactory response $\bm{y}$ to the instruction $\bm{x}$. Therefore, a positive judgment $\bm{j}$ is conferred to praise the commendable performance. The response $\bm{y}$ is aligned with the instruction $\bm{x}$ as well as the combined input $[\bm{x},\bm{j}]$.
Align-N: The LLM makes some mistakes in its generation, resulting in an unsatisfactory response $\bm{y}$. Consequently, a negative judgment $\bm{j}$ details the corresponding critiques. For Align-N, $\bm{y}$ is not aligned with original instruction $\bm{x}$. However, when considering $\bm{x}$ and $\bm{j}$ as a whole, $\bm{y}$ is indeed aligned with the combined input $[\bm{x},\bm{j}]$.
Misalign: The authentic negative judgment in Align-N is substituted with a fake positive judgment $\bm{j}$. In this case, the response $\bm{y}$ is not aligned with either the original instruction $\bm{x}$ or the combination of instruction and judgment $[\bm{x},\bm{j}]$.
@@ 4.2 Learning from Contrasting
Figure 2: Generation probability of identical output text under Align-N (left) and Misalign (right) contexts.
Align-N vs. Misalign: The major difference between these two is that they show opposite polarities in the task of $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$. Thanks to the strong in-context learning capabilities of LLMs, the alignment flip from Align-N (aligned) to Misalign (misaligned) is often accompanied by decreased generation probabilities of the response, particularly for tokens that exhibit a strong correlation with the authentic negative judgment. Figure 2 presents a simple example, wherein the response commits a minor capitalization issue. The LLM assigns a considerably higher probability for “a" when taking the authentic negative judgment $\bm{j}^{-}$ instead of the fake positive judgment $\bm{j}^{+}$ as additional input, precisely at the point where the LLM commits the error.
To take advantage of the above contrast, we feed Align-N and Misalign examples to the LLM to get token generation probabilities $p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})$ and $p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{+})$ separately. We consider the tokens that display a substantially increased generation probability when conditioned on $\bm{j}^{-}$ compared to $\bm{j}^{+}$ as inappropriate tokens (e.g., “a” in Figure 2). Concretely, the following criterion is adopted:
$U=$$\{t\;|\;p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})-\lambda\cdot p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{+})>0\}$ (5)
where $\lambda$ is a hyperparameter to tradeoff the precision and recall of detecting inappropriate tokens.
We apply the UT on the identified inappropriate tokens for pushing the LLM to explore alternative generations. Motivated by the focal loss (Lin et al., 2017), we introduce a dynamic weighting mechanism. This mechanism is designed to modulate the penalty applied to inappropriate tokens in proportion to their degree of inappropriateness. For other tokens, we use the standard MLE loss:
$L_{1}=$ $-\frac{1}{N}(\sum_{t\notin U}\log p(y_{t}|\bm{y}_{<t},\bm{x})$ (6)
$+\sum_{t\in U}\alpha p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})^{\gamma}\log(1-p(y_{t}|\bm{y}_{<t},\bm{x})))$
where $\alpha p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})^{\gamma}$ is the dynamic weight term. $\alpha$ and $\gamma$ are two hyper-parameters. A higher value of $p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})$ suggests that the response tokens have a stronger correlation with negative judgments. Consequently, such tokens are more prone to be inappropriate and are thus subjected to a larger unlikelihood penalty.
Align-P vs. Align-N: Despite both Align-P and Align-N are aligned in terms of $[\bm{x},\bm{j}]\xrightarrow{}\bm{y}$, only Align-P is aligned when solely considering the instruction ($\bm{x}\xrightarrow{}\bm{y}$). Essentially, it suggests that the LLM should output different responses depending on whether a negative judgment is incorporated or not. Therefore, the comparison provides valuable information for the LLM to discern satisfactory and unsatisfactory responses. Specifically, we train on this comparison with the following MLE objective:
[…]
@@
Finally, the overall loss of CUT combines the losses from the two contrasts: $L_{\text{CUT}}=L_{1}+L_{2}$.
@@ 4.3 Relation to Prior Solutions
Forward Prediction hopes that the judgment generation could indirectly boost its response generation abilities. In contrast, CUT directly utilizes judgments to teach the LLM how to generate satisfactory responses and avoid unsatisfactory ones.
ILF assumes judgments can always elicit improved responses. This solution essentially learns from such pseudo improved response. Conversely, CUT can directly learn from misaligned data.
Hindsight learns to generate responses of different qualities at the risk of increasing the likelihood of unsatisfactory responses. In comparison to Hindsight, CUT mitigates this issue by incorporating UT objectives for inappropriate tokens.
@@ 5 Experiments
Method Objective ARC HellaSwag MMLU TruthfulQA Avg. AlpacaEval
LLaMA2 Base - 59.72 81.39 54.97 36.28 58.09 1.87
Forward Prediction MLE 56.91 81.03 54.35 34.28 56.64 7.11
Mistake Analysis MLE 58.53 80.62 54.20 38.80 58.04 12.55
Hindsight MLE 58.11 81.33 55.33 35.61 57.60 10.22
ILF MLE 58.36 81.15 53.76 37.03 57.58 4.01
CDF DPO 58.79 81.07 55.48 41.84 59.3 3.11
CUT (ours) MLE+UT 60.84 81.44 55.78 49.33 61.85 61.06
LLaMA2-chat Base - 58.02 79.89 54.52 45.44 59.47 81.09
Forward Prediction MLE 52.22 78.16 53.06 37.69 55.28 33.21
Mistake Analysis MLE 58.36 78.00 53.77 43.19 58.33 79.29
Hindsight MLE 53.92 78.58 54.15 39.01 56.42 36.67
ILF MLE 58.36 81.15 53.76 45.65 59.73 79.31
CDF DPO 58.81 80.04 54.98 51.51 61.34 83.22
CUT (ours) MLE+UT 58.45 79.86 55.00 52.58 61.47 90.73
Table 2: Results on General Instruction-following. Objective column denotes the fine-tuning objective.
Model rouge1 rouge2 rougeL rougeLsum
LLaMA2 Base 12.91 6.33 10.10 10.87
Forward Prediction 42.42 28.02 38.45 38.51
Mistake Analysis 27.45 13.90 22.84 23.03
Hindsight 38.33 25.49 35.26 35.29
ILF 28.51 16.68 25.36 25.44
CDF 11.31 7.45 10.23 10.77
CUT (ours) 45.39 28.40 39.84 39.89
LLaMA2-chat Base 29.21 15.00 22.78 23.44
Forward Prediction 42.44 28.12 38.48 38.46
Mistake Analysis 40.40 26.11 35.00 35.15
Hindsight 41.02 27.48 37.42 37.46
ILF 39.21 27.93 34.35 34.66
CDF 33.90 19.81 28.01 28.18
CUT (ours) 45.31 29.04 39.96 40.12
Table 3: Results on the summarization task.
We implement CUT in two alignment settings, namely, online alignment and offline alignment, to demonstrate the overall effectiveness of CUT. Subsequently, to highlight the immense potential of aligning LLMs with judgments, we establish a comparison between learning from rewards and learning from judgments. The details of the model implementations are provided in Appendix A.1.
@@ 5.1 Offline Alignment
The offline setting utilizes off-the-shelf instruction-response-judgment triplets for alignment. This aims to check the feasibility of the CUT in learning from judgments prior to initiating the costly process of model-specific judgment annotation.
Tasks. We experiment on a general instruction-following task, and a specific NLP task (summarization). For Instruction following, we train models with 1317 examples from Shepherd (Wang et al., 2023a) and evaluate models on four ranking-based and one generation-based LLM benchmarks. The ranking benchmarks are 25-shot ARC, 10-shot HellaSwag, 5-shot MMLU, and 0-shot TruthfulQA. The generation benchmark is AlpacaEval, where GPT4 judges the winning rate of the responses generated by our models against DaVinci003. For Summarization, we use the dataset with judgment annotations from Saunders et al. (2022). We train models on 10827 training examples and report ROUGE scores (Lin, 2004) on 1939 test examples. See Appendix A.3 for more details.
Baselines. We experiment with two base models, LLaMA2-13b and LLaMA2-chat-13b, to demonstrate the efficacy of CUT in both cold-start and warm-start scenarios, respectively. The baselines include the base model without further fine-tuning, and the three groups of judgment-based alignment methods: 1 The base form of Forward Prediction described in Eq. 1 (Weston, 2016) and Mistake Analysis that uses self-generated judgments and unguided fine-tuning strategy (Chen et al., 2024); (2) The base form of Hindsight described in Eq. 4 (Zhang et al., 2023); (3) The base form of ILF described in Eq. 3 (Scheurer et al., 2022) and CDF (Yu et al., 2023) that change the learning objective from MLE to Direct Preference Optimization (DPO) (Rafailov et al., 2023).
@@
Results. The results of the general instruction-following and summarization are presented in Table 2 and 3, respectively. For cold-start scenarios (LLaMA2 as the base model), CUT improves the winning rate on AlpacaEval from 1.87 to 61.06, where CUT beats the 175B DaVinci003 and surpasses the best baseline (Mistake Analysis) by 48.51 points. Moreover, CUT improves the base model by 13.08 points on TruthfulQA. This implies that CUT can effectively mitigate hallucinations. Conversely, most baselines improve marginally or experience performance drops on TruthfulQA. This is likely due to their application of the MLE objective on error-prone responses, which reduces factuality in response generation. In terms of ARC, HellaSwag, and MMLU, CUT remains competitive with the base model, indicating CUT suffers less from the alignment tax problem (Ouyang et al., 2022). For single NLP task (i.e., summarization) experiments, CUT surpasses the best baseline (i.e., Forward Prediction) by 1.38 rougeLsum scores. Overall, the results show that CUT is effective in transforming LLMs into both performant generalist and specialist models.
@@
The performance improvements of warm-start scenarios (LLaMA2-chat as the base model) are consistent with the cold-start ones, showcasing the efficacy of CUT in learning from judgments in both cold-start and warm-start scenarios. The two ILF methods (ILF and CDF) outperform methods from Forward Prediction and Hindsight groups on AlpacaEval in warm-start but perform worse in cold-start scenarios. This may be due to that ILF methods heavily relies on the base model in producing high-quality improved responses, making it less effective in cold-start scenarios.
Method ARC HeSw MMLU TQA AlpacaEval
Different Model Size
LLaMA2-7b-chat 51.45 78.63 43.60 43.71 71.40
+ CUT 53.16 79.23 46.95 51.40 86.94
LLaMA2-13b-chat 58.02 79.89 54.52 45.44 81.09
+ CUT 58.45 79.86 55.00 52.58 90.73
LLaMA2-70b-chat 65.27 83.89 63.07 53.09 92.70
+ CUT 66.30 84.00 62.71 55.45 93.04
Different Backbone Models
Mistral-7b-it-v1 53.67 74.00 54.66 55.54 69.33
+ CUT 54.27 75.70 54.98 57.61 82.75
gemma-7b-it 51.37 71.67 51.66 30.35 78.51
+ CUT 51.96 72.53 52.28 30.72 81.96
llama3-8b-it 62.20 78.83 65.81 51.65 93.79
+ CUT 61.83 78.91 65.60 51.82 94.09
Table 4: Effect of CUT on different model sizes and different instruction-tuned models. HeSw denotes HellaSwag and TQA denotes TruthfulQA.
Adaptability of CUT. Table 4 presents the impact of CUT framework on a diverse array of models, spanning across multiple model sizes and various instruction-tuned backbone architectures. This examination enables a multifaceted understanding of CUT’s effectiveness and its potential scalability across different model configurations. The upper part of Table 4 focuses on the model sizes, which are analyzed on the LLaMA2-chat family across three distinct scales: 7B, 13B, and 70B. CUT consistently improves the performance across all sizes of the LLaMA2-chat models. This shows that CUT could be scaled up into larger models. Progressing beyond model sizes, the bottom part of Table 4 broadens the scope to include various instruction-tuned backbone models - Mistral-7b-instruct-v1 (Jiang et al., 2023), gemma-7b-it (Team et al., 2024), and llama3-8b-instruct [^1] https://llama.meta.com/llama3. CUT consistently elevates performance across almost all evaluated tasks. This exploration extends the effectiveness of CUT beyond a single model family, shedding light on its adaptability and utility across different model architectures.
Model Generalist Specialist
@@
CUT 52.58 40.12
- $L_{1}$ 39.01 37.46
- first part of $L_{2}$ 26.37
- second part of $L_{2}$ 47.24 33.92
- Inappropriate Token Detection 0 0
- Dynamic Weighting 48.84 40.05
Table 5: Effect of CUT designs. We report the results on TruthfulQA (Acc.) and summarization test set (rougeLsum) for general instruction-following (Generalist) and Summarization (Specialist) respectively. “-” indicates no Align-P examples in the Generalist training set.
Framework Design. To investigate the effectiveness of two contrasts employed by CUT, we perform ablation studies by eliminating certain training signals. The results are shown in Table 5. Removing the contrast between Align-N and Misalign (- $L_{1}$) substantially reduces the performance of TruthfulQA. This finding highlights that the UT objective plays a crucial role in mitigating hallucinations. The exclusion of the contrast between Align-P and Align-N can be implemented in two ways. We can either remove the first part or the second part of $L_{2}$. As seen, the impact of removing Align-P is more pronounced than removing Align-N on the summarization task. This may be attributed to the necessity of positive examples for adapting the LLM to a specific task. Furthermore, we introduce an additional ablated variant in which the inappropriate token detection (Eq. 5) is omitted (- Inappropriate Token Detection). Concretely, we simply apply UT for all tokens in misaligned responses instead. Intriguingly, we find that this approach fails to converge during training. This observation underscores the importance of inappropriate token detection. Lastly, removing the dynamic weighting term ($p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})^{\gamma}$ in Eq. 6) also impacts the effectiveness of the CUT framework, particularly in general instruction-following tasks.
Model #J ARC HellaSwag MMLU TruthfulQA AlpacaEval
LLaMA2-chat - 58.02 79.89 54.52 45.44 81.09
CUT (ours) 1317 58.45 79.86 55.00 52.58 90.73
CUT 1+ (online iteration-1) 1000 58.02 79.55 54.62 50.56 89.75
CUT 2+ (online iteration-2) 1000 57.94 79.18 54.83 51.67 90.23
CUT 3+ (online iteration-3) 1000 58.11 79.99 55.00 52.69 91.04
CUT 4+ (online iteration-4) 1000 58.36 79.23 55.02 52.56 91.68
CUT 5+ (online iteration-5) 1000 58.45 79.19 55.20 52.96 90.68
Table 6: The results of online iterative alignment. #J denotes the number of judgment data used in each iteration.